OpenStudy (anonymous):

Determine the average atomic mass of helium, which has two isotopes having atomic masses 4.0026 amu and 3.0160 amu. The relative abundance of the two isotopes is 99.9999 percent and 0.0001 percent, respectively.

OpenStudy (whpalmer4):

So the average atomic mass is computed with a weighted average. If you have atomic masses \(A,B,C\) with relative abundances \(a\%, b\%, c\%\), the average atomic mass is \[\frac{a}{100}*A + \frac{b}{100}*B+\frac{c}{100}*C\] you only have two isotopes, so you can just ignore the \(c,C\) part.